solve by any method you prefer
x^2 = 11 - x

Answer by stanbon(48535)

x^2 = 11 - x
x^2+x-11=0
Using the quadratic formula get the following:
x=[-1+sqrt(1+44)]/2 or x=[-1-sqrt(45)]/2
x=[-1+3sqrt5]2 or x=[-1-3sqrt5]/2
Cheers,
Stan H.